OpenStudy (anonymous):

y=x-4 x/ 1 / 2 / 3 y/ / /

OpenStudy (calculusxy):

substitute all the values of x to where x is in the equation. \[y = x - 4\] e.g. \[y = 1 - 4\]

OpenStudy (anonymous):

thank you !